Blanchard Company manufactures a single product that sells for $160 per unit and whose total variable costs are $128 per unit. The company’s annual fixed costs are $464,000.

(1) Prepare a contribution margin income statement for Blanchard Company at the break-even point.

(2) Assume the company’s fixed costs increase by $130,000. What amount of sales (in dollars) is needed to break even?

BLANCHARD COMPANY
Contribution Margin Income Statement (at Break-Even)
Amount Percentage of sales
0%
$0
Break-Even Point in Dollars
Choose Numerator: / Choose Denominator: = Break-Even Point in Dollars
/ = Break-even point in dollars
0
Answer #1
BLANCHARD COMPANY
Contribution Margin Income Statement (at Break-Even)
Amount Percentage of sales
Sales (14500*160) 2320000 100%
Variable cost 1856000 80%
Contribution margin 464000 20%
Fixed cost 464000
Operating income $0
Break-Even Point in Dollars
Choose Numerator: / Choose Denominator: = Break-Even Point in Dollars
Fixed cost / contribution margin ratio = Break-even point in dollars
464000 / 0.20 = 2320000

Contribution margin ratio = (160-128)/160 = 20%